# A Survey of Cross-domain Graph Learning: Progress and Future Directions

### v1

@@ IV-C Discussions
Compared with structure-oriented approaches, focusing on learning cross-domain feature commonalities to build graph models also demonstrates its own distinct advantages in handling CDGL especially when the feature information between the source domains and target domains differs significantly. However, because feature-oriented approaches often skip the dedicated extraction of structural information, they may encounter conflicts between structural and feature patterns when training a cross-domain graph model, limiting the performance upper bound.
@@ IV-D Significance and Future Implications
Figure 5: The core ideas of mixture-oriented approaches. a) Feature-Structure Mixture: Feature information is prioritized, and structure information is integrated afterward. b) Structure-Feature Mixture: Structure information is prioritized, and feature information is integrated afterward. c) GNN-based Unified Mixture: Structural and feature information are jointly extracted and processed using a unified extractor during GNN training. d) Flatten-based Unified Mixture: Graph data is transformed into textual sequences through a flattening process, enabling large language models (LLMs) to handle both structural and feature information for downstream tasks. This categorization details two distinct integration philosophies: Sequential-Mixture (a-b) and Unified-Mixture (c-d).
@@ V Mixture-oriented Approach
Mixture-oriented approaches are capable of performing cross-domain learning on both structural and feature information simultaneously, exploring the commonalities of structure and feature either individually or in pairs. This has been the most mainstream research direction in CDGL. Based on the different ways of combining structure and feature information, we categorize Mixture-oriented approaches into two types: Sequential Mixture and Unified Mixture.
@@ V-A Sequential Mixture
Similar to sequential ensemble learning techniques [89, 90], the Sequential Mixture approach typically processes structural information and feature information in a sequential manner. This can involve first processing the structural information, followed by the introduction of feature information. For instance, SDGA [44] first leverages a graph convolutional network (GCN) to extract high-order structural information, followed by an adversarial transformation module with shift parameters to align the source graph feature distribution with that of the target domain, thereby achieving sequential mixture. TAPE [45] leverages LLMs to generate textual explanations as enriched features, which are then used by GNNs to integrate structural information, ultimately enhancing the performance of node classification on text-attributed graphs. GSPT [47] proposes to construct context using random walks first, which is used to represent structural information of the graph, and then feed the context into a transformer to perform masked feature reconstruction, which means the feature information is reconstructed afterward. SOGA [46] first leverages the structural properties of the target graph through Structure Consistency optimization before enhancing feature information through Information Maximization, enabling effective adaptation to the target domain without access to source data. Alternatively, the method can prioritize the feature information before incorporating the structural information. An example of this is UniGraph [49] starts with textual features (raw text features associated with nodes) being fed to the language model encoder (DeBERTa, E5). After processing the textual features, the resulting embeddings are then passed to the GNN encoder for further propagation across the graph. PTGB [43] prioritizes feature transformation and alignment through a data-driven atlas mapping approach before capturing structural properties during GNN pre-training. ARC [48] aligns features across different graph datasets into a common anomaly-sensitive space before utilizing residual-based graph encoding to capture high-order structural information for anomaly detection. AnyGraph [50] employs a feature-centric pre-processing approach using SVD and feature mapping before handling graph structures through an expert routing mechanism and high-order connectivity encoding, allowing for effective cross-domain adaptation. GraphLoRA [51] leverages a structure-aware contrastive low-rank adaptation module that integrates structural information from the target graph, mitigating structural discrepancies, while its Structure-aware Maximum Mean Discrepancy (SMMD) aligns node feature distributions across graphs, ensuring both structural and feature-level adaptation. SAMGPT [52] first aligns and adapts features across domains, then learns universal structure tokens during multi-domain pre-training, and finally utilizes holistic and specific prompts to transfer both unified and domain-specific structural knowledge to downstream tasks. MDGFM [53] first projects domain-specific features into a unified semantic space using unique domain tokens, and then aligns graph topologies through a graph topology-aware alignment mechanism that, via contrastive learning, extracts universal structure tokens for effective prompt-based adaptation.
@@ V-B Unified Mixture
In contrast to the Sequential Mixture approach, the Unified Mixture approach typically considers structural and feature information from a unified perspective during the training process, with the goal of achieving an organic mixture. As shown in Figure 5, unified mixture approaches typically involve two steps: (1) employing a unify function $Unify(\cdot)$ to extract graph structures and features simultaneously to generate node/edge/graph-level representations $H_{rep}$, and (2) a task head function $TaskH(\cdot)$ is then leveraged to mapping the representations and output the predicted results, as illustrated below:
Mixture Unifying: $H_{rep}=Unify(\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{X}),$ (5)
Prediction: $\hat{Y}=TaskH(H_{rep}),$ (6)
where $\mathcal{V},\mathcal{E},\mathcal{X}$ denotes the set of nodes, edges, and node features, respectively. $\mathcal{A}$ is the adjacency matrix of the given graph.
@@
Traditionally, the strategies for $Unify(\cdot)$ and $TaskH(\cdot)$ are part of a cohesive framework. The $Unify(\cdot)$ strategy is typically built upon the message-passing mechanism, where the differences lie in the ways of passing or the applied information encoding techniques. Subsequently, $TaskH(\cdot)$ is realized through a GNN-based task head, such as fully connected layers, which then outputs the task prediction results based on the $H_{rep}$ from $Unify(\cdot)$. However, with the advent of the powerful representation and reasoning capabilities of LLMs in NLP, flatten-based unifying strategies have been introduced into graph learning. These strategies can fundamentally transform graph data into semantically rich node or token sequences. LLMs then leverage these sequences to perform reasoning and generate corresponding task results [91]. To help comprehensively understand these unified mixture approaches, the following sections will illustrate these two sets of unify-task strategies in detail.
GNN-based Unified
@@
In general, GNN-based unified mixture approaches usually adopt similar selection strategies of task head functions while differing according to the passing ways or the applied information encoding techniques. For example, DA-GCN [54], UDA-GCN [54], COMMANDER [56], CCDR [61], GAST [62], DH-GAT [63], ALEX [64], ACT [65], ALCDR [69], and DGASN [67] all adopt domain adaptation training frameworks, differing mainly in their approaches to unifying data from source and target domains. Specifically, DA-GCN and UDA-GCN utilize graph convolutional networks (GCNs) with domain adaptation, which explores the initial possibilities for Unified Mixture CDGL by employing GCNs to aggregate structural and feature information together in a preliminary unified manner. To better handle the unified structure and feature information, more advanced methods have been introduced, such as graph attention mechanisms, hyperbolic representations, and contrastive learning. COMMANDER employs a graph-attentive encoder for domain alignment, providing more flexibility and dynamic focus during feature aggregation. CCDR proposes a unified contrastive learning framework for cross-domain recommendation, utilizing both intra-domain and inter-domain contrastive learning to enhance representation learning and effectively transfer user preferences between domains. GAST uses a hybrid graph attention mechanism along with a POS-Transformer to integrate syntactic and sequential semantics, enhancing both structural and semantic representation. DH-GAT applies hyperbolic graph attention networks to model hierarchical structures effectively, leveraging the power of hyperbolic space to represent latent relationships. ALEX leverages singular value decomposition (SVD) and domain discriminators to unify cross-domain representations, facilitating effective feature alignment. ACT employs contrastive learning with domain alignment for anomaly detection, focusing on learning similarities and differences between domains to generate cohesive representations. ALCDR uses an optimal transport-based anchor link learning mechanism to adaptively transfer knowledge across domains, providing a robust means of cross-domain unification. DGASN employs a graph attention network (GAT) to simultaneously learn node and edge embeddings, and using adversarial domain adaptation to align the source and target distributions, thus achieving domain-invariant graph representations.
@@
Additionally, CDGEncoder [58], PGPRec [70], STGP [68], GCOPE [21], UniGLM [73], CrossLink [71], and MDGPT [72] all involve a pretraining phase, but they differ in their information encoding or propagation methods during pretraining. Specifically, the CDGEncoder uses a multi-view GNN encoder combined with attention mechanisms to aggregate contextual and topological views. PGPRec introduces personalized graph prompts and uses contrastive learning during pretraining to enhance user/item representations. STGP utilizes a two-stage prompting strategy to adapt both domain and task-specific properties across spatio-temporal tasks. GCOPE employs graph coordinators to align diverse graph datasets during pretraining, mitigating negative transfer. UniGLM considers introducing LLMs as the encoder for text attributes in the traditional pre-training paradigm with a tailored contrastive learning approach based on the adaptive positive sample selection to align features and structures. CrossLink leverages a decode-only transformer to model the temporal evolution of dynamic graphs for link prediction. Finally, MDGPT uses domain tokens and dual prompts to unify and adapt multi-domain graph features, enhancing cross-domain generalization.
@@
In addition to the above two mainstream paradigms, domain generalization, multi-task learning, and federated learning have also been explored. For example, CrossHG-Meta [59] proposes to process structural and feature information simultaneously by aggregating heterogeneous information from multiple semantic contexts and then designs a unified domain generalization framework for CDGL, especially for heterogeneous graphs. METABrainC [60] proposes a pre-training cross-domain framework via meta-learning techniques using multiple self-supervised tasks on source datasets and adapts it to target tasks via fine-tuning, enhancing cross-domain brain connectome analysis. GCFL [57] employs a clustering algorithm to group graph models trained on clients based on parameter similarities, thereby achieving a unified representation of structural and feature information within a federated learning framework, ultimately enhancing cross-domain generalization.
@@
Referring to the input of LLMs, flatten-based Unified Mixture approaches focus on converting graphs into textual descriptions by transforming them into sequences of nodes or tokens. These approaches then develop tailored learning strategies to allow LLMs to process these sequences and generate task-specific outputs effectively. For instance, GIMLET [74] introduces a unified graph-text approach called Generalized Position Embedding, which transforms graphs into token sequences and seamlessly combines graph structure with LLM-driven text processing. GraphTranslator [75] employs a Translator module to convert node embeddings into token sequences, enabling effective integration of pre-trained graph models and LLMs, allowing LLMs to handle both pre-defined and open-ended graph tasks. GraphGPT [76] leverages a dual-stage instruction tuning mechanism to align graph structural and semantic information with LLMs, enhancing cross-domain generalization through both graph structural alignment and node feature transfer. GITA [77] combines graph visualization with textual conversion, using a Vision-Language Model (VLM) to jointly reason over visual and textual graph data, providing a comprehensive integration of graph semantics and structure. InstructGraph [78] uses a structured format verbalizer to convert graph data into text sequences, combined with instruction tuning and preference optimization to improve graph reasoning and reduce hallucinations, making LLMs suitable for graph-centric tasks. TEA-GLM [79] employs a linear projection module to align pre-trained Graph Neural Network (GNN) representations with LLM token embeddings, enabling zero-shot predictions across multiple graph tasks without modifying the LLM itself. GOFA [80] introduces a hybrid architecture that interleaves GNN layers with a pre-trained LLM, combining structural and semantic modeling abilities while transforming graph data into flattened text descriptions for LLM-based text generation. HiGPT [81] utilizes a heterogeneous graph tokenizer to transform complex graph structures into textual representations, employing a novel instruction-tuning framework to enhance LLM’s understanding of intra- and inter-node relationships. GraphCLIP [82] uses a graph-summary contrastive learning approach by generating textual descriptions of subgraphs with an LLM, which improves zero-shot and few-shot transfer capabilities through graph prompt tuning and invariant learning. GraphWiz [83] leverages the GraphInstruct dataset to train LLMs in solving graph computational problems, using an instruction-following framework that transforms graphs into textual sequences, allowing explicit reasoning paths for improved task performance.

### v2

@@ IV-C Discussions
Compared with structure-oriented approaches, focusing on learning cross-domain feature commonalities to build graph models also demonstrates its own distinct advantages in handling CDGL, especially when the feature information between the source domains and target domains differs significantly. However, since feature-oriented approaches often skip the dedicated extraction of structural information, they may encounter conflicts between structural and feature patterns when training a cross-domain graph model, limiting the performance upper bound.
@@ IV-D Significance and Future Implications
Figure 5: Mixture-oriented CDGL. (a) Feature-Structure Mixture: first extracts transferable features, then integrates structural knowledge. (b) Structure-Feature Mixture: first extracts shared structural patterns, then incorporates feature information. (c) GNN-based Unified Mixture: a unified extractor jointly learns cross-domain structural and feature patterns. (d) Flatten-based Unified Mixture: graphs are flattened into token sequences, leveraging the powerful reasoning ability of LLMs to align and extract shared cross-domain knowledge. (a)-(b) follow a sequential-mixture integration paradigm, whereas (c)-(d) adopt a unified-mixture strategy. The red flame icon denotes trainable modules.
@@ V Mixture-oriented Approach
Mixture-oriented approaches are capable of performing cross-domain learning on both structural and feature information simultaneously, exploring the commonalities of structure and feature either individually or in pairs. This has been the most mainstream research direction in CDGL. Based on the different ways of fusing structure and feature information, we classify Mixture-oriented approaches into two types: Sequential Mixture and Unified Mixture.
@@ V-A Sequential Mixture
Similar to sequential ensemble learning techniques [rayana2016sequentialensemble1, yang2020sequentialensemble2], the Sequential Mixture approach typically processes structural information and feature information in a sequential manner. As shown in Figure 5, sequential mixture approaches typically involve three steps: (1) employing a feature extractor $FExtract(\cdot)$ to extract shared graph features, mapping $\mathcal{X}_{S}\in\mathbb{D}_{G_{S}}$ into $\mathcal{X}^{{}^{\prime}}_{S}$, (2) then, leveraging a structure extractor $SExtract(\cdot)$ to extract shared graph structures, mapping $\mathcal{A}_{S}\in\mathbb{D}_{G_{S}}$ into $\mathcal{A}^{{}^{\prime}}_{S}$, and (3) finally, a graph model $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is trained on the new source domains $\mathbb{D}^{{}^{\prime}}_{G_{S}}$ derived by $FExtract(\cdot)$ and $SExtract(\cdot)$ and subsequently generalized to the target domains $\mathbb{D}_{T}$ via a task-specific head function $TaskH(\cdot)$, which output the predicted results.
Feature Extract: $\mathcal{X}^{{}^{\prime}}_{S}=FExtract(\mathcal{X}_{S}),$ (20)
Structure Extract: $\mathcal{A}^{{}^{\prime}}_{S}=SExtract(\mathcal{A}_{S}),$ (21)
Generalization: $\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{{}^{\prime}}_{T},\mathcal{X}^{{}^{\prime}}_{T},\mathcal{A}^{{}^{\prime}}_{T})),$ (22)
Here, the order of the feature extractor and the structure extractor can be swapped depending on the design.
For instance, SDGA [qiao2023SDGA] first leverages GCN to extract high-order structural information, followed by an adversarial transformation module with shift parameters to align the source graph feature semantic distribution with the target domain, thereby transferring structures and features in a sequential manner. GSPT [song2024GSPT] first samples node sequences via Node2Vec [grover2016node2vec] to capture cross-domain structural contexts, then maps node attributes into a unified space using a pretrained LLM (e.g., SentenceBERT [reimers2019sentenceBert]). It pre-trains a Transformer through masked feature reconstruction on these sequences to mine feature commonalities, and finally transfers to target domains without tuning, similar to  [liu2024OFA]. SOGA [mao2024SOGA] aims to adapt a GNN pre-trained on source domains to target domains by leveraging the Structure Consistency and Information Maximization objectives. Specifically, the two objectives are computed separately for structures and features, thereby fine-tuning the pre-trained GNN to capture the structural and feature commonalities between the source and target domains, achieving cross-domain transfer. MLDGG [tian2025mldgg] sequentially learns domain-generalizable knowledge by first employing a structure learner to refine noisy edges and capture cross-domain structural patterns, and then applying a representation learner to disentangle domain-invariant semantics from domain-specific variations in node embeddings, achieving meta-level generalization to unseen domains.
Alternatively, the method can also prioritize the feature information before incorporating the structural information. An example of this is UniGraph [he2024unigraph]. Specifically, it first leverages a pretrained LM [he2020deberta, wang2022textE5] to unify node feature spaces on source domains. Then, it utilizes a GNN encoder to encode the structural information and introduces PPR-sampled subgraphs and a masked‐autoencoder learning objective to pretrain the LM and GNN encoder, distilling cross-domain commonalities. Finally, it can be effectively transferred to target domains via in-context learning or graph instruction tuning on an LLM [touvron2023llama, touvron2023llama2]. PTGB [yang2023ptgb] first designs a data‐driven atlas mapping pipeline to project and align node features from multiple source domains into a common low-dimensional space, then pre-trains a GNN via a novel two-level contrastive learning objective to mine structural commonalities, and finally fine-tunes the pretrained GNN on target domains for cross-domain transfer. ARC [liu2024ARC] first applies a smoothness-based feature alignment module to project and sort node attributes from source domains into a shared, anomaly-sensitive feature space; it then uses an ego-neighbor residual graph encoder to extract high-order structural anomaly patterns; finally, it performs anomaly detection on target domains via a cross-attentive in-context scoring module without any retraining. AnyGraph [xia2024anygraph] first applies SVD to align the node-feature dimensions of multiple source domains into a shared embedding space. It then introduces a mixture-of-experts architecture—each specializing in a distinct structural pattern—and uses a lightweight routing mechanism to dynamically select the best expert for each graph. Finally, selected experts combined with corresponding feature encoders are jointly pre-trained on source domains to capture transferable cross-domain commonalities for target domains. GraphLoRA [yang2024graphlora] first injects a low-rank adaptation network alongside the frozen pre-trained GNN and then employs a structure-aware Maximum Mean Discrepancy (SMMD) loss to align node feature distributions between source and target graphs; it then uses structure-aware contrastive learning to infuse structural knowledge from the target graph into the model, ensuring effective cross-graph transfer at both structural and feature levels. SAMGPT [yu2025SAMGPT] first aligns and adapts features across domains, then learns universal structure tokens during multi-domain pre-training, and finally utilizes holistic and specific prompts to transfer both unified and domain-specific structural knowledge to target domains. MDGFM [wang2025MDGFM] first projects each source domain’s features into a shared tokenized semantic space to capture cross-domain semantic commonalities, then introduces the topology-aware contrastive learning strategy to align the graph structures on source domains, and finally transfers the learned tokens and GNN encoder to target domains via prompting.
@@ V-B Unified Mixture
In contrast to the Sequential Mixture approach, the Unified Mixture approach typically considers structural and feature information from a unified perspective during the training process, with the goal of achieving an organic mixture. As shown in Figure 5, unified mixture approaches typically involve two steps: (1) employing a unify function $Unify(\cdot)$ to extract graph structures and features simultaneously, mapping $\mathcal{A}_{S}\in\mathbb{D}_{G_{S}}$ and $\mathcal{X}_{S}\in\mathbb{D}_{G_{S}}$ into $\mathcal{A}^{{}^{\prime}}_{S}$ and $\mathcal{X}^{{}^{\prime}}_{S}$, and (2) then, a graph model $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is trained on the new source domains $\mathbb{D}^{{}^{\prime}}_{G_{S}}$ derived by $Unify(\cdot)$ and subsequently generalized to the target domains $\mathbb{D}_{T}$ via a task-specific head function $TaskH(\cdot)$, which output the predicted results.
Mixture Unifying: $\mathcal{X}^{{}^{\prime}}_{S},\mathcal{A}^{{}^{\prime}}_{S}=Unify(\mathcal{X}_{S},\mathcal{A}_{S}),$ (23)
Prediction: $\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{{}^{\prime}}_{T},\mathcal{X}^{{}^{\prime}}_{T},\mathcal{A}^{{}^{\prime}}_{T})),$ (24)
@@
Traditionally, the strategies for $Unify(\cdot)$ and $TaskH(\cdot)$ are part of a cohesive framework. The $Unify(\cdot)$ strategy is typically built upon the message-passing mechanism, where the differences lie in the ways of passing or the applied information encoding techniques. $TaskH(\cdot)$ is typically implemented using GNN-based heads (e.g., MLPs), which take the representations from $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ and output predictions for the target domains $\mathbb{D}_{T}$. However, with the advent of the powerful representation and reasoning capabilities of LLMs in NLP, flatten-based unifying strategies have been introduced into graph learning. These strategies can fundamentally transform graph data into semantically rich node or token sequences. LLMs then leverage these sequences to perform reasoning and generate corresponding task results [ijcai2024GFMSurvey]. To help comprehensively understand these unified mixture approaches, the following sections will illustrate these two sets of unified-task strategies in detail.
Graph-based Unified
@@
In general, GNN-based unified mixture approaches usually adopt similar selection strategies of task head functions while differing according to the message passing ways or the information encoding techniques. For example, DA-GCN [Guo2021DAGCN], UDA-GCN [wu2020UDAGCN], COMMANDER [ding2021COMMANDER], CCDR [xie2022CCDR], GAST [zhang2022GAST], DH-GAT [xu2023DHGAT], ALEX [yuan2023alex], ACT [wang2023ACT], ALCDR [zhao2023ALCDR], DGASN [shen2023DGASN], CMPGNN [wang2025bridgingCMPGNN], and AEGOT-CDKE [wu2025crossAEGOTCDKT] all adopt domain adaptation training frameworks, differing mainly in their approaches to unifying data from source and target domains. Specifically, DA-GCN [Guo2021DAGCN] and UDA-GCN [wu2020UDAGCN] utilize graph convolutional networks (GCNs) [kipf2016gcn] with domain adaptation, which explores the initial possibilities for Unified Mixture CDGL by employing GCNs to aggregate structural and feature information together in a preliminary unified manner. To better handle the unified structure and feature information, more advanced methods have been introduced, such as graph attention mechanisms [velivckovic2018gat], hyperbolic representations, and contrastive learning. COMMANDER [ding2021COMMANDER] employs a graph-attentive encoder for domain alignment, providing more flexibility and dynamic focus during feature aggregation. CCDR [xie2022CCDR] proposes a unified contrastive learning framework for cross-domain recommendation, utilizing both intra-domain and inter-domain contrastive learning to enhance representation learning and effectively transfer user preferences between domains. GAST [zhang2022GAST] uses a hybrid graph attention mechanism along with a POS-Transformer to integrate syntactic and sequential semantics, enhancing both structural and semantic representation. DH-GAT [xu2023DHGAT] applies hyperbolic graph attention networks to model hierarchical structures effectively, leveraging the power of hyperbolic space to represent latent relationships. ALEX [yuan2023alex] leverages SVD [stewart1993SVD] and domain discriminators to unify cross-domain representations, facilitating effective feature alignment. ACT [wang2023ACT] employs contrastive learning with domain alignment for anomaly detection, focusing on learning similarities and differences between domains to generate cohesive representations. ALCDR [zhao2023ALCDR] uses an optimal transport-based anchor link learning mechanism to adaptively transfer knowledge across domains, providing a robust means of cross-domain unification. DGASN [shen2023DGASN] employs a graph attention network [velivckovic2018gat] to simultaneously learn node and edge embeddings, and uses adversarial domain adaptation to align the source and target distributions, thus achieving domain-invariant graph representations. CMPGNN [wang2025bridgingCMPGNN] inserts cross-domain edges via a link predictor to enable message passing between domains, unifying structure and feature spaces while using a mutual information loss to preserve discriminativeness and mitigate domain distribution shifts. AEGOT-CDKT [wu2025crossAEGOTCDKT] achieves cross-domain transfer by aligning node embeddings from source and target domains using Graph Optimal Transport (GOT), reducing domain discrepancies, and enhancing knowledge tracing, modeled as node classification.
@@
Additionally, CDGEncoder [hassani2022CDGEncoder], PGPRec [yi2023PGPRec], STGP [hu2024stgp], GCOPE [zhao2024GCOPE], UniGLM [fang2024uniglm], CrossLink [huang2024CrossLink], MDGPT [yu2024MDGPT], MDP-GNN [lin2025MDPGNN], UniGraph2 [he2025unigraph2], and GIT [wang2025GIT] all involve a pretraining phase, but they differ in their information encoding or propagation methods during pretraining. Specifically, CDGEncoder [hassani2022CDGEncoder] uses a multi-view GNN encoder combined with attention mechanisms to aggregate contextual and topological views. PGPRec [yi2023PGPRec] introduces personalized graph prompts and uses contrastive learning during pretraining to enhance user/item representations. STGP [hu2024stgp] utilizes a two-stage prompting strategy to adapt both domain and task-specific properties across spatio-temporal tasks. GCOPE [zhao2024GCOPE] employs graph coordinators to align diverse graph datasets during pretraining, mitigating negative transfer. UniGLM [fang2024uniglm] considers introducing LLMs as the encoder for text attributes in the traditional pre-training paradigm with a tailored contrastive learning approach based on the adaptive positive sample selection to align features and structures. CrossLink [huang2024CrossLink] leverages a decode-only transformer to model the temporal evolution of dynamic graphs for link prediction. Finally, MDGPT [yu2024MDGPT] uses domain tokens and dual prompts to unify and adapt multi-domain graph features, enhancing cross-domain generalization. MDP-GNN [lin2025MDPGNN], a cross-domain pre-training framework, first learns a feature converter for each domain to map node attributes into a shared semantic space, then injects a set of shared pivot nodes whose learned links blend graph structures across domains, and finally minimizes a Wasserstein-distance term to align their overall distributions, where the pre-trained GNN model can be effectively fine-tuned or prompted in cross-domain downstream tasks. UniGraph2 [he2025unigraph2] focuses on cross-domain multimodal graph learning. It introduces a mixture-of-experts module [wang2023MoEs] to align and fuse cross-domain features, and trains a GNN with a unified self-supervised loss that reconstructs node attributes as well as graph structures, thus preserving rich semantic and structural information. The same pipeline also works for single-modality graphs by simply disabling the unused modality branch. Compared with GFT [wang2024gft], which performs cross-domain learning by directly converting each whole graph into a computation tree (a structure generation CDGL approach), GIT [wang2025GIT] first builds a task tree—adding a virtual task node and linking only task-relevant parts of the graph—then pre-trains a GNN on many such task trees from diverse domains with a contrastive reconstruction loss [zhu2020GRACEContrastive], optionally instruction-tunes the encoder with a handful of labeled task trees from the target domain, and finally solves downstream tasks by transforming each new graph into its task tree and fine-tuning the pre-trained GNN model plus a lightweight classifier. OMOG [liu2024OMOG] first leverages a frozen LLM (i.e., Sentence-BERT)[reimers2019sentenceBert] to unify node text features and applies multi-hop SGC[wu2019SGC] to encode structural information from cross-domain graphs. It then pre-trains a dedicated expert model for each source graph to capture domain-specific knowledge. By computing the adaptability between these expert models and input graphs from new domains, OMOG enables effective few-shot and zero-shot learning on downstream tasks.
@@
In addition to the above two mainstream paradigms, domain generalization, multi-task learning, and federated learning have also been explored. For example, CrossHG-Meta [zhang2022CrossHGMeta] proposes to process structural and feature information simultaneously by aggregating heterogeneous information from multiple semantic contexts and then designs a unified domain generalization framework for CDGL, especially for heterogeneous graphs. METABrainC [yang2022METABrainC] proposes a pre-training cross-domain framework via meta-learning techniques using multiple self-supervised tasks on source datasets and adapts it to target tasks via fine-tuning, enhancing cross-domain brain connectome analysis. GCFL [xie2021gcfl] employs a clustering algorithm to group graph models trained on clients based on parameter similarities, thereby achieving a unified representation of structural and feature information within a federated learning framework, ultimately enhancing cross-domain generalization.
@@
Referring to the input of LLMs, flatten-based Unified Mixture approaches focus on converting graphs into textual descriptions by transforming them into sequences of nodes or tokens. These approaches then develop tailored learning strategies to allow LLMs to process these sequences and generate task-specific outputs effectively. For instance, GIMLET [zhao2023gimlet] introduces a unified graph-text approach called Generalized Position Embedding, which transforms graphs into token sequences and seamlessly combines graph structure with LLM-driven text processing. GraphTranslator [zhang2024graphtranslator] employs a Translator module to convert node embeddings into token sequences, enabling effective integration of pre-trained graph models and LLMs, allowing LLMs to handle both pre-defined and open-ended graph tasks. GraphGPT [tang2024graphgpt] leverages a dual-stage instruction tuning mechanism to align graph structural and semantic information with LLMs, enhancing cross-domain generalization through both graph structural alignment and node feature transfer. GITA [wei2024gita] combines graph visualization with textual conversion, using a Vision-Language Model (VLM) to jointly reason over visual and textual graph data, providing a comprehensive integration of graph semantics and structure. InstructGraph [wang2024instructgraph] uses a structured format verbalizer to convert graph data into text sequences, combined with instruction tuning and preference optimization to improve graph reasoning and reduce hallucinations, making LLMs suitable for graph-centric tasks. TEA-GLM [wang2024TeaGLM] employs a linear projection module to align pre-trained Graph Neural Network (GNN) representations with LLM token embeddings, enabling zero-shot predictions across multiple graph tasks without modifying the LLM itself. GOFA [kong2025gofa] introduces a hybrid architecture that interleaves GNN layers with a pre-trained LLM, combining structural and semantic modeling abilities while transforming graph data into flattened text descriptions for LLM-based text generation. HiGPT [tang2024higpt] utilizes a heterogeneous graph tokenizer to transform complex graph structures into textual representations, employing a novel instruction-tuning framework to enhance LLM’s understanding of intra- and inter-node relationships. GraphCLIP [zhu2024graphclip] uses a graph-summary contrastive learning approach by generating textual descriptions of subgraphs with an LLM, which improves zero-shot and few-shot transfer capabilities through graph prompt tuning and invariant learning. GraphWiz [chen2024graphwiz] leverages the GraphInstruct dataset to train LLMs in solving graph computational problems, using an instruction-following framework that transforms graphs into textual sequences, allowing explicit reasoning paths for improved task performance.